Point charge q1 = -6.00 * 10-6 C is on the x-axis at x = -0.200 m. Point charge q2 is




on the x-axis at x = +0.400 m. Point charge q3 = +3.00 * 10-6 C is at the origin. What




is q2 (magnitude and sign) (a) if the net force on q3 is 6.00 N in the +x direction; (b)




if the net force on q3 is 6.00 N in the -x direction?

1
Expert's answer
2022-01-03T10:33:33-0500

Part (a)


"F=\frac{k(q_1q_2)}{r^2}"


"F_{net}=F_{13}+F_{23}"


"6=8.988×10^9[\frac{(-6×10^{-6})(3×10^{-6})}{(-0.2)^2}+\frac{(q_2)(3×10^{-6})}{(0.4)^2}]"


"q_2=+5.96×10^{-5}\>C"


Part (b)


"-6=8.988×10^9[\frac{(-6×10^{-6})(3×10^{-6})}{(-0.2)^2}+\frac{(q_2)(3×10^{-6})}{(0.4)^2}]"


"q_2=-1.16×10^{-5}\>C"